# A Survey of Backdoor Attacks and Defenses on Large Language Models: Implications for Security Measures

### v3

@@ 6 Discussion and Open Challenges
forked edges, for tree= grow’=0, draw, rounded corners, align=center, parent anchor=east, child anchor=west, l sep=20pt, s sep=15pt, font = , anchor=center, [Open
Challenges [Backdoor
Attack [Trigger Design
Section 6.1[ How to design more covert and universal backdoor
attack triggers?]] [Clean-labe
Section 6.2[How to deploy clean-label backdoor attacks on
generation tasks?]] [Without Fine-tuning
Section 6.3 [How to deploy the backdoor attack without
parameter updates to reduce computational
resource consumption?]] ] [Defend Against
Backdoor Attack [Defenses
Section 6.4 [How to design a high-performance and universal
defense algorithm?]] [Backdoor Evaluation
Section 6.5 [How to proactively check if a model has
been compromised?]] ] [Others
Section 6.6 [Interpretation Analysis; Evaluation Metrics; Uniform Benchmark]] ]
@@ 6.1 Trigger Design
Existing backdoor attacks demonstrate promising results on victim models. However, the deployment of backdoor attacks often requires embedding triggers in samples, which may compromise the fluency of those samples. Importantly, samples containing triggers have the potential to alter the original semantics of the instances. Additionally, the insertion of explicit triggers considerably increases the risk of the backdoor being detected by defense algorithms, such as in scenarios involving instruction poisoning (Wang et al., 2023a, ) and ICL poisoning (Zhao et al., 2024b, ). Hence, the design of more covert and universal triggers still needs to be considered.
@@ 6.4 General and Effective Defenses
Defending against backdoor attacks is crucial for safeguarding the application of large language models. Although existing defense algorithms can achieve the expected outcomes, their generality remains limited. For instance, the ONION (Qi et al., 2021a, ) algorithm can effectively defend against character-level trigger backdoor attacks but fails to counter sentence-level trigger backdoor attacks (Chen et al., 2021b, ). Furthermore, current defense algorithms rely on additional training steps or multiple iterations of search to identify and mitigate backdoor threats. This not only has the potential to consume substantial computational resources but also necessitates further enhancements in efficiency. Consequently, given the intricacy and diversity of backdoor attacks, the development of versatile and high-performance defense algorithms represents a crucial research imperative.
@@ 6.5 Backdoor Evaluation
At present, language models are in a passive defensive stance when confronted with backdoor attacks, lacking efficacious methodologies to determine whether they have been compromised by the implantation of backdoors. For instance, Zhao et al., 2024a propose a new defense algorithm based on the assumption that the model had been compromised through weight poisoning. Although previous research has demonstrated good defensive outcomes, these are predicated on the assumption that the language model has been compromised. Indiscriminate defense not only consumes resources but also has the potential to impair the performance of unaffected models. Considering the insufficiency of current evaluation methods, designing a lightweight yet effective assessment method is a problem worthy of investigation.
@@ 6.6 Others
Interpretation Analysis It is noteworthy that due to the inherent black-box nature of neural networks, backdoor attacks are challenging to interpret. Investigating the interpretability of backdoor attacks is crucial for devising more efficient defense algorithms. Comprehending the mechanisms behind backdoor attacks can better expose their internal characteristics, providing essential insights for the development of defense strategies.
Evaluation Metrics In settings with a limited sample label space, the attack success rate is commonly used as an evaluation metric. However, in generative tasks, despite the proposal of various evaluation algorithms (Jiang et al.,, 2023), a unified standard of assessment is still lacking. Furthermore, evaluating the stealthiness of backdoor attacks is also a worthy topic of discussion.
Uniform Benchmark The establishment of uniform benchmarks is crucial for assessing the effectiveness of backdoor attacks and defense algorithms, necessitating standardized poisoning ratios, datasets, baseline models, and evaluation metrics.

### v5

@@ 6 Discussion and Open Challenges
forked edges, for tree= grow’=0, draw, rounded corners, align=center, parent anchor=east, child anchor=west, l sep=40pt, s sep=15pt, font = , anchor=center, [Open
Challenges , fill=orange!5 [Backdoor
Attack, fill=yellow!30 [Trigger Design
Section 6.1 , fill=green!10 [ How to design more covert and universal backdoor
attack triggers? , fill=orange!25]] [Clean-Labe
Section 6.2 , fill=green!10 [How to deploy clean-label backdoor attacks on generation
tasks? , fill=orange!25]] [Without Fine-tuning
Section 6.3 , fill=green!10 [How to deploy the backdoor attack without parameter
updates to reduce computational resource
consumption?, fill=orange!25]] ] [Defend Against
Backdoor Attack, fill=yellow!30 [Defenses
Section 6.4 , fill=green!10 [How to design a high-performance and universal
defense algorithm?, fill=orange!25]] [Backdoor Evaluation
Section 6.5 , fill=green!10 [How to proactively check if a model has been
compromised?, fill=orange!25]] ] [Others
Section 6.6 , fill=yellow!30 [Interpretation Analysis; Evaluation Metrics;
Uniform Benchmark , fill=green!10]] ]
@@ 6.1 Trigger Design
Existing backdoor attacks demonstrate promising results on victim models. However, the deployment of backdoor attacks often requires embedding triggers in samples, which may compromise the fluency of those samples. Importantly, samples containing triggers have the potential to alter the original semantics of the instances. Additionally, the insertion of explicit triggers considerably increases the risk of the backdoor being detected by defense algorithms, such as in scenarios involving instruction poisoning (Wang et al., 2023a) and ICL poisoning (Zhao et al., 2024c). Hence, the design of more covert and universal triggers still needs to be considered.
@@ 6.4 General and Effective Defenses
Defending against backdoor attacks is crucial for safeguarding the application of large language models. Although existing defense algorithms can achieve the expected outcomes, their generality remains limited. For instance, the ONION (Qi et al., 2021a) algorithm can effectively defend against character-level trigger backdoor attacks but fails to counter sentence-level trigger backdoor attacks (Chen et al., 2021b). Furthermore, current defense algorithms rely on additional training steps or multiple iterations of search to identify and mitigate backdoor threats. This not only has the potential to consume substantial computational resources but also necessitates further enhancements in efficiency. Consequently, given the intricacy and diversity of backdoor attacks, the development of versatile and high-performance defense algorithms represents a crucial research imperative.
@@ 6.5 Backdoor Evaluation
At present, language models are in a passive defensive stance when confronted with backdoor attacks, lacking efficacious methodologies to determine whether they have been compromised by the implantation of backdoors. For instance, Zhao et al. (2024b) propose a new defense algorithm based on the assumption that the model had been compromised through weight poisoning. Although previous research has demonstrated good defensive outcomes, these are predicated on the assumption that the language model has been compromised. Indiscriminate defense not only consumes resources but also has the potential to impair the performance of unaffected models. Considering the insufficiency of current evaluation methods, designing a lightweight yet effective assessment method is a problem worthy of investigation.
@@ 6.6 Others
Interpretation Analysis: It is noteworthy that due to the inherent black-box nature of neural networks, backdoor attacks are challenging to interpret. Investigating the interpretability of backdoor attacks is crucial for devising more efficient defense algorithms. Comprehending the mechanisms behind backdoor attacks can better expose their internal characteristics, providing essential insights for the development of defense strategies.
Evaluation Metrics: In settings with a limited sample label space, the attack success rate is commonly used as an evaluation metric. However, in generative tasks, despite the proposal of various evaluation algorithms (Jiang et al., 2023), a unified standard of assessment is still lacking. Furthermore, evaluating the stealthiness of backdoor attacks is also a worthy topic of discussion.
Uniform Benchmark: The establishment of uniform benchmarks is crucial for assessing the effectiveness of backdoor attacks and defense algorithms, necessitating standardized poisoning ratios, datasets, baseline models, and evaluation metrics.